OpenStudy (anonymous):

Please check my answer! write the terms of the series and find the sum. \[\sum_{k=1}^{4} (k+9)^{2}\] N =4 534

OpenStudy (anonymous):

just expand the series 100, 121....

OpenStudy (asnaseer):

yes - your answer is correct @marcybaby :)

OpenStudy (asnaseer):

wait - what you just wote does not agree with what you wrote in the question

OpenStudy (anonymous):

woooohooo!! I'm killing myself over here!

OpenStudy (anonymous):

how u got that terms?

OpenStudy (asnaseer):

in your question you wrote an answer of 534 - correct?

OpenStudy (anonymous):

but what about the terms?

OpenStudy (asnaseer):

how did you get that answer?

OpenStudy (anonymous):

I'm gonna need a new brain after this!

OpenStudy (asnaseer):

no need to panic :)

OpenStudy (asnaseer):

just explain how you got your answer of 534

OpenStudy (anonymous):

1/6 (541 N + 57 N^2 + 2 N^3)

OpenStudy (asnaseer):

where did that expression come from?

OpenStudy (anonymous):

HA!!! i have NO idea!

OpenStudy (anonymous):

I dont know how the hell I got that answer anymore!

OpenStudy (asnaseer):

ok, lets take it one step at a time. you are told that the k'th term is equa to \((k+9)^2\). so, given that, what do you think will be the first first (i.e. when k=1) ?

OpenStudy (asnaseer):

*first term

OpenStudy (anonymous):

i'm so lost now! :(

OpenStudy (asnaseer):

ok, the sum is given as:\[\sum_{k=1}^{4} (k+9)^{2}\]agreed?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

the \(\sum\) symbol means "sum this expression" and the "k=1" at the bottom of this and "4" at its top mean that "k" should take on the values from 1 to 4. agreed?

OpenStudy (anonymous):

yes.

OpenStudy (asnaseer):

good - this is basically a short hand way of writing the sum. so, if we expand the sum, we will get:\[\sum_{k=1}^{4} (k+9)^{2}=(1+9)^2+(2+9)^2+(3+9)^2+(4+9)^2\]can you see how that follows?

OpenStudy (anonymous):

I do!

OpenStudy (asnaseer):

gr8! so all you need to do now is work out the numbers on the right and bingo! you should get to your answer. :)

OpenStudy (anonymous):

so it shoule be (10)^2? or (1)^2&(9)^2?

OpenStudy (asnaseer):

you evaluate the expression inside the braces first, e.g.:\[(1+9)^2=(10)^2=100\]

OpenStudy (anonymous):

okay! 100, 121, 144, 169

OpenStudy (asnaseer):

perfect! now just add these up...

OpenStudy (anonymous):

534

OpenStudy (asnaseer):

thats it! great work! :)

OpenStudy (asnaseer):

see - no new brain required :)

OpenStudy (anonymous):

duh! LOL i did that you know. but i thought i had it all wrong

OpenStudy (anonymous):

thank you so much!!!! i need a drink!

OpenStudy (asnaseer):

learn to trust yourself :)

OpenStudy (asnaseer):

and you're welcome :)

OpenStudy (anonymous):

xoxo

OpenStudy (asnaseer):

go on - you deserve that drink now...

OpenStudy (anonymous):

I wish! i have more to do.

OpenStudy (asnaseer):

hehe - no worries, I'm sure the drinks will still be there when you've finished! :)

OpenStudy (anonymous):

Ha! thanks a million! first drink is dedicated to you! :)

OpenStudy (asnaseer):

cheers to that! :)